OpenStudy (anonymous):

is the y-value of sin(x) when x= -360° is y is 0?

OpenStudy (johnweldon1993):

Well we know that 360 would be equivalent to \(\large 2\pi\) it doesn't matter what direction we travel around the unit circle for that since we will always end up back at \(\large 2\pi\) \[\large sin(2\pi) = sin(-2\pi) = 0\]